# Forward modeling fluctuations in the DESI LRGs target sample using image simulations

### v1

@@ 5.2 Source injection strategy
To fully maximize computing resources, without sacrificing the authenticity of our simulated sources, we develop a method to place the sources optimally. We put sources in a hexagonal pattern on each brick. For a given minimum distance between sources, this method achieves the maximum number density. The distance between the closest sources is 360 pixels, or 94". We also eliminate sources that touch pixels containing bright galaxies, bright stars, and clusters. Pixels containing these sources are masked out in the imaging systematics tests (see Section 6). Further, such areas often correspond to a large blob that contains many sources and is computationally expensive to process. Given the areas will not later be used for analysis, the Obiwan pipeline is made more efficient by ignoring them. The Legacypipe code produces a bitmask map for each brick; we exclude areas that are near optically bright and “medium-bright” stars, WISE $W1$-bright stars, large galaxies, and globular clusters. These correspond to the “maskbits” data product bits 1, 11, 8, 12, and 13, respectively.
@@ 6 Obiwan with DR9 images
We inject galaxies in the DECaLS North Galactic Cap (NGC) and South Galactic Cap (SGC) regions, and the footprint is shown in Figure 7. If no pixel is masked, 100 galaxies are injected for each brick. This corresponds to a distance of 360 pixels between simulated galaxies, which is 94 arcsec. We deliberately chose this large separation to avoid blending on WISE images. In total, 6.6 million galaxies are simulated in the SGC, and 8.5 million galaxies are simulated in the NGC. The density of final Obiwan LRGs is about 30% that of DESI LRGs. Figure 8 shows the time consumed for each brick. Obiwan runs on Permutter nodes at NERSC. With 128 threads on a single node, one node can process 8 bricks in parallel. On average it takes 11 minutes to finish one brick with 16 threads. Despite the fast speed of each brick, the pipeline is still inefficient. The problem is that the parallelization is limited by the memory (RAM) of the CPU available during the simulation. To read all the images needed for processing, we can only process a maximum of 8 bricks at a time. This means that with 16 threads available for computation, only 1 thread is active most of the time. The parallelization happens in the source fitting stage. As Obiwan fits many fewer sources than a normal reduction run, the efficiency on a per-source basis is therefore lower compared to the full image processing.
@@
6.1 Measurement bias
Overall, the flux measurement is negatively biased. Figure 9 shows the measurement flux bias as a function of input magnitude in each band. The dashed line in the middle is the median value in each flux bin. It shows that the bias is consistent across all magnitude bins, and it exists in all optical and infrared bands. To understand whether these features are a result of pipeline structures, or some features in the images, we designed 4 parallel runs: On the first run, we first generate synthetic Gaussian background noise that is much smaller than a typical Legacy Survey image. We inject galaxies into these images. In this setting, there are very few galaxies in the image, and no blending would occur here. The image is also almost noiseless. The second run is similar to the first run. The difference is that we set the Gaussian noise to the level of a typical Legacy Survey Image. In the third setting, we take the images on the second run and add the real galaxies back with their model parameters recorded in the DR9 data release. With this setting, we have a fully simulated image without any contamination. Blending may occur in such images, but it is free of error from background subtraction, cosmic ray contamination, etc. In the final setting, we use the real images and add simulated galaxies to these images, which is the same way as our previous production runs. The 4 production runs use the same input galaxy setting on the same set of images. In total, 442 bricks are processed for this comparison. Table 1 shows the results of these production runs. The results show that flux bias does not appear on these non-real images. This means that there is no algorithmic negative bias in the pipeline. The bias from real images is caused by other effects that can not be easily simulated. Fortunately, the bias is small enough that it has little effect on how we select galaxies as LRGs. If we correct the bias in our simulated outputs as a function of flux:
[…]
Figure 9: The density (blue) plot of output flux minus input flux. The black dashed line is the 84th, 50th, and 16th percentile for a given input magnitude. The red dashed line is centered at 0 on the y-axis. The difference between the red line and the 50th-percentile line shows that the flux measurement is consistently biased at all input flux levels in all bands. The four plots show results in $g,r,z,W1$ band respectively.
The population of galaxies that pass the target selection cuts with the new flux values only changes by 0.77%. Therefore this bias has a negligible effect when it comes to determining the target selection of LRGs.
[…]
Table 1: Median flux difference (output-input) for 4 production runs. In the first column, simulated galaxies are added to images containing tiny Gaussian noise (standard deviation of $10^{-4}$ nanomaggies per pixel). In the second column, simulated galaxies are added to images with Gaussian noise expected in the weight image. In the third column, galaxies are added to images with regular level gaussian noise, plus idealized model galaxy images for the DR9 catalog. In the fourth column, simulated galaxies are added to real DR9 images, as in our main Obiwan runs. In the real images, the fluxes are consistently biased low. This does not happen in other image sets. For the "negligible noise" set, there is approximately no bias present. In the "regular noise" and "Noise + model galaxies" sets, the flux values are biased high.
@@ 6.2 Morphology test
Morphologically, there are 6 types of light profiles included in Legacypipe, and 5 of them are used in our analysis. The un-used one is the "DUP" type, a confirmed star by external surveys. Below is a list of the 5 types of profiles; each type needs a different number of parameters to describe its shape.
@@
Figure 11: A matrix showing the proportion of sources measured given the light profile denoted on the $y$-axis given the input profile used to simulate the source, denoted on the $x$-axis. The fitting pipeline penalizes model complexity and thus will use a more complex model to measure flux only if the image signal to noise is high enough to statistically detect a preference. See the text for details on the specific profile and statistical thresholds.
Figure 10 shows the proportion of different types of galaxies in DESI LRGs and Obiwan LRGs. We see good agreement between these two. This is a result of the debiasing method described in Section 5.1, as galaxy types are closely related to galaxy shapes. For example, PSF and REX sources have an ellipticity of 0, so we need the sample to have the right proportion of sources with an ellipticity of 0. With a precise estimate of morphological distribution from the input truth data, we obtain an accurate shape distribution in the output Obiwan LRGs. Figure 11 shows the scatter between the input and recovered output morphological type of our simulated galaxies. Most of the galaxies are recovered as their input type, or a simpler type (meaning a model of fewer parameters). As is shown in equation 5.1, the model used depends on several terms. The first term decides how much the image deviates from the model, and it is controlled by the noise level in that location. The third term does not depend on the variance in the image, but only on the number of parameters used in the model. When the image is more noisy, the first term becomes less important, and the pipeline favors a model with fewer parameters. Additionally, there is minimal misclassification of exponential galaxies (EXP) as de Vaucouleurs (DEV) galaxies, or vice versa. This could be useful for identifying LRGs with star-forming disks.
@@ 6.3 Imaging Systematics
We use the term ‘Imaging Systematics’ as a phrase to mean false fluctuations in the target galaxy density arising from varying observing conditions and astrophysical foreground in imaging surveys. Measurements of galaxy properties are contaminated by numerous effects such as Galactic extinction, denoted as $E(B-V)$; the local density of stars at some flux threshold, denoted as stardens; the detection limit of a point source, denoted as psfdepth; and the amount by which the Earth’s atmosphere disperses the light we gather with our telescopes, denoted as psfsize. For foreground contaminations like stardens and $E(B-V)$, we use maps  [^7] We specifically used https://data.desi.lbl.gov/public/ets/target/catalogs/dr9/1.1.1/pixweight/main/resolve/dark/pixweight-1-dark.fits. produced as described in section 4.5.2 of [52]. The maps are produced using HealPix [63] at a resolution $N_{\rm side}=256$. For effects related to observing conditions, like psfsize and psfdepth, we use direct outputs from the catalog produced by the image reduction pipeline. We see a slight improvement in the matching between Obiwan LRGs and DESI LRGs, particularly in using the catalog level psfdepth, compared with using pixel values in HealPix. psfdepth depends on the list of images processed in a certain footprint. The combination of image coadds varies at a resolution smaller than HealPix resolution, so it is also conceptually favorable to test the trends of psfdepth on a catalog level.
We attach the values of stardens ($E(B-V)$ is already in the catalog) from the HealPix map to the catalogs of DESI LRGs and Obiwan LRGs. We use Obiwan input as randoms. Using the Obiwan input as randoms is equivalent to using a general random catalog. It has the advantage of reducing Poisson noise when tested with Obiwan LRGs. The data counts are binned into eight even percentiles of the given property. Then, the reference random sample is binned in the same way. We divide the data counts by the reference and normalize them by the total number of reference counts to data (so that the overall mean is 1). We obtain results separately for data in the NGC, shown in Fig. 12, and SGC, shown in Fig. 13.
To quantify the consistency between the data and the simulated trends, via a $\chi^{2}$ statistic, we first need covariance matrixes for DESI LRGs and Obiwan LRGs. We use:
@@
With a degree of freedom of 8, a trend that fully resembles DESI LRGs trend should have a $\chi^{2}<13.4$ at 90% confidence level, or $\chi^{2}<15.5$ at 95% confidence level. In the NGC, the plot shows that there is no clear systematic trend in psfsize g, psfsize r, psfsize z, and psfdepth W1, either predicted or observed. We see that the Obiwan LRGs successfully predict the trends observed in psfdepth g, psfepth r, and psfdepth z. The depth of the images indicates the noise level of galaxies, which indicates that the systematic trends are from photometric scattering. However, we do see trends in extinction and stellar density that are not predicted by Obiwan. We investigate these further in Section 7.2.
In the SGC, we see qualitative agreement in the slopes of the trends, but the $\chi^{2}$ results suggest a poor match between the DESI LRGs trends and those that Obiwan predicts. For psfdepth, the mismatch is especially obvious in high-depth regions. We explain the reason for this mismatch in Section 7.1. In short, this mismatch is not a result of failure in Obiwan simulation. It arises from not having a perfect truth input. The mismatch can be corrected if we can match the color distribution of DESI LRGs with Obiwan LRGs.
Figure 12: Predicted and observed trends between LRG density and imaging properties of the North Galactic Cap. The shaded curves are the results for DESI LRGs, while the blue lines are for synthetic Obiwan LRGs. From top-to-bottom and left-to-right, the first 4 plots are psfdepth trend in $g,r,z,W1$ bands, respectively. The next three plots show the trends against psfsize in the $g,r,z$ bands, respectively. Next is the trend against Galactic extinction as measured by [65], and finally, the trend against stellar density. The observed trends with psfdepth are predicted. There is no obvious trend with psfsize in either sample. The trends in $E(B-V)$ and stellar density are not predicted by the synthetic Obiwan-LRGs, and we believe this is due to contamination in the $E(B-V)$ map (see Section 7.2).
@@ 7 Testing the Obiwan procedure
There are several trends between the observed LRG density and imaging properties that we were not able to simulate in the fiducial Obiwan settings. In this section, we discuss components not simulated by Obiwan, which causes a mismatch in imaging systematics trends that we see in figure 12 and 13. Section 7.1 discusses the dependency of imaging systematics on intrinsic galaxy brightness, and how imperfections in the "truth catalog" cause mismatch by not having the perfect magnitude distribution. Section 7.2 discusses uncertainties in the $E(B-V)$ maps and its correlation with the number density of DESI LRGs. Section 7.3 discusses signs of un-simulated PSF error and that it could potentially change the characteristics of the LRG population near the color selection boundaries. Section 7.4 discusses other non-simulated secondary effects that could bias our results, but we do not see a clear sign of their impact. Section 7.5 discusses potential issues with the approach of using survey property maps for imaging systematics mitigation, which is the most widely used method for current galaxy surveys. We identify two factors that are not commonly considered with this method, and how forward modeling tools like Obiwan can help.
@@ 7.1 Correlation With Galaxy Brightness
The trends with psfdepth888psfdepth is the depth for a point source and galdepth is the depth for a fiducial galaxy. They are similar, and here we use psfdepth. The conclusions apply to both. still have a small discrepancy between what is observed and predicted, especially at the greatest depths. This originates from the small mismatch in the magnitude distribution between DESI LRGs and Obiwan LRGs, shown in Fig. 14. The panels show that Obiwan LRGs are generally fainter than DESI LRGs, especially in the $z$-band. We investigate possible explanations for this mismatch.
@@
One possible reason is that the COSMOS Deep data is not strictly noiseless. Although overall COSMOS Deep is a clean sample in optical bands, the flux W1 in the WISE band does not have deep photometry, as is shown in figure 3. The luminosity cut shown in equation 4.4 is an important source of photometric scattering and it involves utilizing the WISE W1 band. We do not find evidence of any strong discrepancies induced by imperfect WISE band input, but we can not rule out the possibility that it could cause a mismatch in color distribution in other bands.
@@
A second reason is that we selected truth galaxies from a small footprint, and the color distribution of galaxies in this small footprint deviates from the overall sample distribution. Fig. 15 validates this possibility. In the left plot, we see LRGs selected from the COSMOS deep catalog have a dip at $z$-band fiber magnitude around 20.5. As we use COSMOS deep catalog as "truth input", this feature gets populated into Obiwan LRGs. Our findings suggest that the magnitude distribution of LRGs in the COSMOS Deep region is not strictly a representative sample of DESI LRGs as a whole. To compensate the loss of LRG density in the brighter $z$-band fiber magnitude range, it also has a higher density LRGs at the fainter part of $z$-band fiber magnitude (>21.4) than the overall magnitude distribution of DESI LRGs. As fiber $z$ magnitude and (total) $z$ magnitude are highly correlated, this causes more Obiwan LRGs to appear in the faint $z$ band magnitude region. Furthermore, the DESI LRGs in the COSMOS deep footprint has a fiber $z$ magnitude distribution very different from the overall DESI LRG distribution in NGC. It suggests that this footprint is too small to average out photometric noise in the shallower DR9 imaging, which is another evidence that this footprint is not big enough for a selection of statistically representative LRG sample. Thus, our discrepancy comes from not having a large enough footprint of the COSMOS Deep catalog. The mismatch does not appear only on $z$-band, but compared with other bands, $z$-band is easier to see the difference by eye. Thus, we choose $z$-band data for illustration. The result here suggests that we should use a larger footprint to generate the “truth catalog” in future work. If everything worked correctly, we would see a better match between DESI LRGs in COSMOS deep (or any other deep data used), and the DESI LRGs in the NGC (or any other larger footprint). We should also be able to see a much smaller difference between the $z$-band magnitude distribution of Obiwan LRGs and DESI LRGs.
Figure 14: Magnitude distribution of DESI LRGs (Orange) and Obiwan LRGs (Blue) in the North Galactic Cap (NGC, solid lines) and the South Galactic Cap (SGC, dashed lines). Overall, the magnitude distribution of Obiwan LRGs agrees well with DESI LRGs. However, Obiwan LRGs has a higher density near the faint end of $z$-band magnitude. This feature is related to how we choose the truth catalog to populate Obiwan sources and it is further explained in Figure 15. The discrepancy here likely contributes to the mismatch in the SGC systematics map seen in Figure 13. The mismatch appears in the region with high depth, which is unique to the SGC. It can be corrected by adding weights to the Obiwan LRGs to make their magnitude distribution match with the DESI LRGs, which is shown in Figure 17.
@@
Figure 15: The $z$-band fiber magnitude and total magnitude distributions of LRGs in different catalogs and regions. The blue histograms show LRGs in the COSMOS Deep catalog. The black histogram shows DESI LRGs cut to the COSMOS Deep footprint. The magenta histograms show DESI LRGs in the whole North Galactic Cap (NGC) region. The red histograms show Obiwan LRGs in the NGC region. The left plot shows a histogram in $z$-band fiber magnitude. The right plot shows the total $z$-band magnitude. The left plot suggests that the COSMOS Deep region is a small region, and DESI LRGs in this same region (Black) have a very different distribution compared to DESI LRGs in a much larger footprint (Magenta). However, LRGs selected from the COSMOS Deep catalog (Blue) show a better resemblance to the histogram of LRGs in a larger footprint (Magenta). The COSMOS Deep LRGs (Blue) have an under-density feature at around $z$-band fiber magnitude $20.5$. This feature gets populated to Obiwan LRGs (Red) so we see the same under-density in Obiwan LRGs in the same $z$-band fiber magnitude range. In the right plot is similar to the $z$-band magnitude plot shown in figure 14, and we re-plot here for better visualization to demonstrate why we see a $z$-band magnitude discrepancy between DESI LRGs and Obiwan LRGs.
@@
So far, we have only looked at trends between the full LRG target sample and the survey property maps. As the strongest trends we predict are with the imaging depth, one could reasonably expect the significance of the trends to depend on the brightness of the LRG target. To investigate this, we split the DESI LRGs sample into two sub-samples: the bright sample ($z$ band magnitude $<20.5$), and the faint sample ($z$ band magnitude $>20.5$). With this split, we observe dramatically different trends, as shown in Fig. 16. For the brighter half of the data, we find no significant trend observed or predicted with the depth. Conversely, a quite strong trend is observed in the faint half of the data, with $\sim$ 15% variation in number density as a function of depth. The trends are well-predicted by Obiwan, with similar performance in both the NGC and SGC. This suggests that a portion of the mismatch found for the SGC in Fig. 13 was due to the excess of faint galaxies in the Obiwan sample predicting a somewhat stronger trend in the full sample than observed in the real data.
@@
The Obiwan LRGs predict the brightness-dependent trends with imaging systematics. However, as is shown in Fig. 14, it is difficult to achieve a perfect match in magnitude distribution between DESI LRGs and Obiwan LRGs. To mitigate this issue, we re-weight Obiwan LRGs so that their magnitude distribution matches with DESI LRGs. We use the basic bin-based reweighting method [^9] https://hsf-training.github.io/analysis-essentials/advanced-python/45DemoReweighting.html to match the magnitude distribution of $g$, $r$, $z$ , $g-r$, and $r-z$ of Obiwan LRGs to DESI LRGs. The matching is successful in that the KS values obtained from the comparison of the Obiwan LRGs and DESI LRGs magnitude distributions are reduced from 0.018, 0.028, 0.038, 0.027, 0.016 to 0.008, 0.008, 0.007, 0.007, 0.006, respectively. We apply these weights to Obiwan LRGs and re-compute the systematics plot with these weights applied.
Figure 17 shows the trends between LRG density and the psfdepth in the SGC, after re-weighting the Obiwan LRGs. Compared to the result in Figure 13 (We repeated relevant plots in figure 17), the $\chi^{2}$ value in Obiwan LRGs is greatly reduced. This process of re-weighting is a rudimentary study of how to improve the matching between real and synthetic sources. In regions of different survey properties, the magnitude distribution varies. This method ignores this variation. In addition, although the re-weighted distribution has a much smaller KS value, it is still not small enough to mathematically declare the distributions of Obiwan LRGs and DESI LRGs are drawn from the same underlying distribution. In our future work, we will address this issue thoroughly.

### v2

@@ 5.2 Source injection strategy
The injected galaxies on the images need to be far from each other while maintaining a large number density. The optimal pattern to achieve this is the hexagonal pattern. This method achieves the maximum number density for a given minimum distance between sources. Thus, we adopt this pattern for source injection. The distance between the closest sources is 360 pixels, or 94". We deliberately chose this large separation to avoid blending on WISE images. We also eliminate sources that touch pixels containing bright galaxies, bright stars, and clusters; the details are in Appendix A.1.
@@ 6 Obiwan with DR9 images
We inject galaxies in the DECaLS North Galactic Cap (NGC) and South Galactic Cap (SGC) regions, and the footprint is shown in Figure 4. If no pixel is masked, 100 galaxies are injected for each brick. In total, 6.6 million galaxies are simulated in the SGC, and 8.5 million galaxies are simulated in the NGC. The density of final Obiwan LRGs is about 30% that of DESI LRGs. The number density of Obiwan LRGs is limited by computational resources, and we put more details in Appendix A.3. We also found that the flux of output galaxies is systematically lower than the injected values. However, this flux bias has a negligible effect on LRG target selection. Appendix D provides an in-depth discussion over this issue.
@@ 6.1 Morphology test
Morphologically, six types of light profiles are included in Legacypipe, and 5 of them are used in our analysis. The unused one is the "DUP" type, a confirmed star by external surveys. Below is a list of the five types of profiles; each type needs a different number of parameters to describe its shape.
@@
Figure 6: A matrix showing the proportion of sources measured given the light profile denoted on the $y$-axis given the input profile used to simulate the source, denoted on the $x$-axis. The fitting pipeline penalizes model complexity and thus will use a more complex model to measure flux only if the image signal to noise is high enough to detect a preference statistically. See the text for details on the specific profile and statistical thresholds.
Figure 5 shows the proportion of different types of galaxies in DESI LRGs and Obiwan LRGs. We see good agreement between these two. This results from the debiasing method described in Appendix C, as galaxy types are closely related to galaxy shapes. For example, PSF and REX sources have an ellipticity of 0, so we need the sample to have the right proportion of sources with an ellipticity of 0. With a precise estimate of morphological distribution from the input truth data, we obtain an accurate shape distribution in the output Obiwan LRGs. Figure 6 shows the scatter between the input and recovered output morphological type of our simulated galaxies. Most galaxies are recovered as their input type, or a simpler type (meaning a model of fewer parameters). As is shown in equation 5.1, the model used depends on several terms. The first term decides how much the image deviates from the model, and it is controlled by the noise level in that location. The third term does not depend on the variance in the image, but only on the number of parameters used in the model. When the image is more noisy, the first term becomes less important, and the pipeline favors a model with fewer parameters. Additionally, there is minimal misclassification of exponential galaxies (EXP) as de Vaucouleurs (DEV) galaxies or vice versa. This could be useful for identifying LRGs with star-forming disks. In Section 7.3, we found a correlation between the morphological type of LRGs and $z$-band psfsize. An accurate shape recovery helps us investigate this issue.
@@ 6.2 Imaging Systematics
[…]
We attach the values of stardens ($E(B-V)$ is already in the catalog) from the HealPix map to the catalogs of DESI LRGs and Obiwan LRGs. We use Obiwan input as randoms. Using the Obiwan input as randoms is equivalent to using a general random catalog. The general random catalog is uniformly distributed in the (RA,DEC) sphere. We also apply the same mask to it as DESI LRGs. The random catalog from Obiwan input has the advantage of reducing Poisson noise when tested with Obiwan LRGs. The data counts are binned into eight even percentiles of the given property. Then, the reference random sample is binned in the same way. We divide the data counts by the reference and normalize them by the total number of reference counts to data (so that the overall mean is 1). We obtain results separately for data in the NGC, shown in Figure 7, and SGC, shown in Figure 8.
To quantify the consistency between the data and the simulated trends via a $\chi^{2}$ statistic, we first need covariance matrixes for DESI LRGs and Obiwan LRGs. We use:
@@
With a degree of freedom of 8, a trend that fully resembles DESI LRGs trend should have a $\chi^{2}<13.4$ at 90% confidence level, or $\chi^{2}<15.5$ at 95% confidence level. In the NGC, the plot shows that there is no clear systematic trend in psfsize g, psfsize r, psfsize z, and psfdepth W1, either predicted or observed. We see that the Obiwan LRGs successfully predict the trends observed in psfdepth g, psfepth r, and psfdepth z. The images’ depth indicates the galaxies’ noise level, which suggests that the systematic trends are from photometric scattering. However, we do see trends in extinction and stellar density that are not predicted by Obiwan. We investigate these further in Section 7.2.
In the SGC, we see qualitative agreement in the slopes of the trends, but the $\chi^{2}$ results suggest a poor match between the DESI LRGs trends and those that Obiwan predicts. The mismatch for psfdepth is pronounced in high-depth regions. We explain the reason for this mismatch in Section 7.1. In short, this mismatch does not result from failure in Obiwan simulation. It arises from not having a perfect truth input. The mismatch can be corrected if we match the color distribution of DESI LRGs with Obiwan LRGs.
Figure 7: Predicted and observed trends between LRG density and imaging properties of the North Galactic Cap. The shaded curves are the results for DESI LRGs, while the blue lines are for synthetic Obiwan LRGs. From top-to-bottom and left-to-right, the first 4 plots are psfdepth trend in $g,r,z,W1$ bands, respectively. The next three plots show the trends against psfsize in the $g,r,z$ bands. Next is the trend against Galactic extinction as measured by [66], and finally, the trend against stellar density. The observed trends with psfdepth are predicted. Neither sample has an obvious trend with psfsize. The trends in $E(B-V)$ and stellar density are not predicted by the synthetic Obiwan-LRGs, and we believe this is due to contamination in the $E(B-V)$ map (see Section 7.2).
@@ 7 Testing the Obiwan procedure
There are several trends between the observed LRG density and imaging properties that we could not simulate in the fiducial Obiwan settings. In this section, we discuss components not simulated by Obiwan, which causes a mismatch in imaging systematics trends that we see in Figure 7 and 8. Section 7.1 discusses the dependency of imaging systematics on intrinsic galaxy brightness, and how imperfections in the "truth catalog" cause mismatch by not having the perfect magnitude distribution. Section 7.2 discusses uncertainties in the $E(B-V)$ maps and its correlation with the number density of DESI LRGs. Section 7.3 discusses signs of un-simulated PSF error and that it could potentially change the characteristics of the LRG population near the color selection boundaries. Section 7.4 discusses other non-simulated secondary effects that could bias our results, but we do not see a clear sign of their impact. Section 7.5 discusses potential issues with using survey property maps for imaging systematics mitigation, the most widely used method for current galaxy surveys. We identify two factors not commonly considered with this method and how forward modeling tools like Obiwan can help. Section 7.6 talks about how the systematics discussed in this work impact cosmological measurements by potentially introducing a varying galaxy bias that is not properly mitigated.
@@ 7.1 Correlation With Galaxy Brightness
The trends with psfdepth still have a small discrepancy between what is observed and predicted, especially at the greatest depths. This originates from the small mismatch in the magnitude distribution between DESI LRGs and Obiwan LRGs, shown in Figure 9. The panels show that Obiwan LRGs are generally fainter than DESI LRGs, especially in the $z$-band. We investigate possible explanations for this mismatch.
@@
One possible reason is that the COSMOS Deep data is not strictly noiseless. Although COSMOS Deep is overall a clean sample in optical bands, the W1 flux in the WISE band does not have deep photometry, as is shown in Figure 19. The luminosity cut shown in equation 4.4 is an important source of photometric scattering, and it involves utilizing the WISE W1 band. We do not find evidence of any strong discrepancies induced by imperfect WISE band input, but we can not rule out the possibility that it could cause a mismatch in color distribution in other bands.
A second reason is the sample variance from the truth catalog. The truth catalog is from a small footprint, so the LRG-like sample selected in this region is slightly different from the LRG-like sample in the full footprint. This sample from the truth catalog serves as seeds that generate Obiwan inputs. We inject these Obiwan inputs to images and further produce Obiwan LRGs. Therefore, the features exclusively of the LRGs from the truth catalog can also be seen in Obiwan LRGs. We demonstrate this effect in Figure 9. In the left plot, we see LRGs selected from the COSMOS deep catalog have a dip at $z$-band fiber magnitude around 20.5. As we use COSMOS deep catalog as "truth input", this feature is also seen in Obiwan LRGs. Due to this sample variance, the COSMOS deep region LRGs also have a higher density LRGs at the fainter part of $z$-band fiber magnitude (>21.4) than the overall magnitude distribution of DESI LRGs.
@@
Thus, our discrepancy comes from not having a large enough footprint of the COSMOS Deep catalog. The mismatch does not appear only on $z$-band, but compared with other bands, $z$-band is easier to see the difference by eye. Thus, we choose $z$-band data for illustration. The result here suggests that we should use a larger footprint to generate the “truth catalog” in future work. If everything worked correctly, we would see a better match between DESI LRGs in COSMOS deep (or any other deep data used), and the DESI LRGs in the NGC (or any other larger footprint). We should also see a much smaller difference between the $z$-band magnitude distribution of Obiwan LRGs and DESI LRGs.
@@
Figure 9: The $z$-band fiber magnitude and total magnitude distributions of LRGs in different catalogs and regions. The blue histograms show LRGs in the COSMOS Deep catalog. The black histogram shows DESI LRGs cut to the COSMOS Deep footprint. The magenta histograms show DESI LRGs in the whole North Galactic Cap (NGC) region. The red histograms show Obiwan LRGs in the NGC region. The left plot shows a histogram in $z$-band fiber magnitude. The right plot shows the total $z$-band magnitude. The left plot suggests that the COSMOS Deep region is a small region, and DESI LRGs in this same region (Black) have a very different distribution compared to DESI LRGs in a much larger footprint (Magenta). However, LRGs selected from the COSMOS Deep catalog (Blue) show a better resemblance to the histogram of LRGs in a larger footprint (Magenta). The COSMOS Deep LRGs (Blue) have an under-density feature at around $z$-band fiber magnitude $20.5$. This feature is transferred to Obiwan LRGs (Red) with Obiwan simulation so we see the same under-density in Obiwan LRGs in the same $z$-band fiber magnitude range.
@@
So far, we have only looked at trends between the full LRG target sample and the survey property maps. As the strongest trends we predict are with the imaging depth, one could reasonably expect the significance of the trends to depend on the brightness of the LRG target. To investigate this, we split the DESI LRGs sample into two sub-samples: the bright sample ($z$ band magnitude $<20.5$), and the faint sample ($z$ band magnitude $>20.5$). With this split, we observe dramatically different trends, as shown in Figure 10.
For the brighter half of the data, we find no significant trend observed or predicted with the depth. Conversely, a quite strong trend is observed in the faint half of the data, with $\sim$15variation in number density as a function of depth. The trends are well-predicted by Obiwan, with similar performance in both the NGC and SGC. This suggests that a portion of the mismatch found for the SGC in Figure 8 was due to the excess of faint galaxies in the Obiwan sample predicting a somewhat stronger trend in the full sample than observed in the real data.
@@
The Obiwan LRGs predict the brightness-dependent trends with imaging systematics. However, as is shown in Figure 9, it is difficult to achieve a perfect match in magnitude distribution between DESI LRGs and Obiwan LRGs. To mitigate this issue, we re-weight Obiwan LRGs so that their magnitude distribution matches with DESI LRGs. We use the basic bin-based reweighting method [^7] https://hsf-training.github.io/analysis-essentials/advanced-python/45DemoReweighting.html to match the magnitude distribution of $g$, $r$, $z$ , $g-r$, and $r-z$ of Obiwan LRGs to DESI LRGs. The matching is successful in that the KS values obtained from the comparison of the Obiwan LRGs and DESI LRGs magnitude distributions are reduced from 0.018, 0.028, 0.038, 0.027, 0.016 to 0.008, 0.008, 0.007, 0.007, 0.006, respectively. We apply these weights to Obiwan LRGs and re-compute the systematics plot with these weights applied.
Figure 11 shows the trends between LRG density and the psfdepth in the SGC, after re-weighting the Obiwan LRGs. Compared to the result in Figure 8 (We repeated relevant plots in Figure 11), the $\chi^{2}$ value in Obiwan LRGs is greatly reduced. This re-weighting process is a rudimentary study of how to improve the matching between real and synthetic sources. In regions of different survey properties, the magnitude distribution varies. This method ignores this variation. In addition, although the re-weighted distribution has a much smaller KS value, it is still not small enough to mathematically declare the distributions of Obiwan LRGs and DESI LRGs are drawn from the same underlying distribution. In our future work, we will address this issue thoroughly.